### v2

# Dynamic Competition between Cooper-Pair and Spin-Density-Wave Condensation

Abstract
Quantum matter phases may co-exist microscopically even when they display competing tendencies. A fundamental question is whether such a competition can be avoided through the elimination of one phase while the other one condenses into the ground state. Here, we present a high-resolution neutron spectroscopy study of the low-energy spin excitations in the high-temperature superconductor La1.855Sr0.145CuO4. In the normal state, we find low-energy magnetic fluctuations at incommensurate reciprocal lattice positions where spin-density-wave order emerges at lower Sr concentration or at high magnetic fields. While these spin excitations are largely suppressed by the emergence of the superconducting spin gap, some low-energy magnetic fluctuations persist deep inside the superconducting state. We interpret this result in terms of a dynamic competition between superconductivity and magnetism, where superconductivity impedes the condensation of low-energy magnetic fluctuations through the formation of magnetically-mediated Cooper pairs.
I Introduction
After more than a hundred years since the discovery of superconductivity, the field still remains a source of fascination for the scientific community. During the last four decades particular interest in the field arose through the discovery of high-temperature superconductivity. Many of these unconventional superconducting condensates emerge either in proximity to or in coexistence with magnetically ordered ground states MathurNAT1998. Yet, superconductivity typically competes with static magnetism (see Fig. 1a) NandiPRL2010; ChangPRB2008. This static competition has been explored by a multitude of techniques monitoring the two order parameters or their onset temperatures as a function of temperature or extrinsic tuning parameters including magnetic field, chemical doping and pressure Grissonnanche14a; huckerPRB2014; blancocanosaPRB2014.
While magnetic long-range order tends to compete with superconductivity, it is widely believed that magnetic fluctuations provide the underlying attractive interaction to establish unconventional Cooper pairing Scalapino2012. This has motivated detailed studies of magnetic excitations in unconventional superconductors using in particular neutron scattering techniques. An experimental signature of superconductivity - found in cuprate, heavy-fermion and iron-based superconductors - is a collective spin-1 excitation below the superconducting transition temperature ROSSATMIGNOD1991; TranquadaPRB2004; ChristensenPRL2004; Stock2008; Mazzone2017; Yu2009. This spectral feature, often referred to as ”resonance“, is thought to arise due to magnetically mediated Cooper pairing as a feedback effect of unconventional superconductivity opening a gap at the Fermi surface. In turn, this is reflected in the spin excitation spectrum by the appearance of a low-energy spin gap, from which the resonance draws its spectral weight ROSSATMIGNOD1991; Dai2001; Mason1993; YamadaPRL1995. In the past, several studies have assessed the interplay between the spin resonance and spin gap on the one hand, and static magnetism on the other one. Such studies have yielded insight into the symmetry of the superconducting order parameter and its effect on the magnetic long-range order Mazzone2017; Dai2015. These studies, however, mainly focused on the effects of static order onto the spin resonance, but did not study how the magnetically mediated Cooper pairs affect the dynamic properties which ultimately lead to magnetic long-range order.
A possible route to address this dynamic competition is through studies focusing on how the formation of Cooper pairs affects the low-lying magnetic fluctuations. Notably, unconventional superconductivity often appears close to second-order magnetic phase transitions, where the symmetry breaking magnetic order parameter emerges through a continuous quantum phase transition that is linked to a magnetic soft mode. If unconventional Cooper pairs are built from low-energy magnetic fluctuations, the formation of Cooper pairs is likely impacting the low-energy modes associated to the magnetic long-range order (see Fig. 1b-d). Thus, insight into this dynamic competition between magnetic fluctuations and unconventional superconductivity can be gained through studies of the low-energy magnetic excitation spectrum. Ideally, such a study is performed either via non-thermal tuning across the static antiferromagnetic (AFM)-superconducting phase boundary ChangPRL2009, or by studying the temperature dependence of the fluctuations in a material without magnetic order close to the AFM phase boundary across (see Fig. 1a). The latter approach requires a material with a large to ensure an appreciable temperature window over which the emergent spin gap below the superconducting resonance can be investigated.
Here we focus on La1.855Sr0.145CuO4 which is a model material that fulfills these requirements. It displays no magnetic long-range order at zero magnetic field, and features a superconducting phase below = 36 K with a spin gap 4 meV ChangPRL2009; ChangPRB2008 appearing at an incommensurate quartet of wavevectors. Spin stripe-order tranquadaEvidenceStripeCorrelations1995 is recovered upon reduction of the Sr content ( 0.135) ChangPRB2008; KofuPRL2009 and when a magnetic field ( 7 T) is a applied along the crystallographic -axis ChangPRL2009. Using time-of-flight neutron spectroscopy with an elastic resolution of = 80 eV full-width at half maximum (FWHM), we observe low-energy magnetic fluctuations for at the wavevectors characteristic of the incipient stripe order. These fluctuations are gradually wiped out by the emergent spin gap, but we observe persistent spectral weight below 1 meV transfer. The result is interpreted as a near-condensation of low-energy spectral weight that is impeded by the opening of the spin gap. Notably, we observe a dynamic competition where superconductivity prevents spin-density-wave ordering.
II Experimental Details
The experiment was carried out using the IN5 time-of-flight neutron spectrometer Ollivier2011 at the Institut Laue-Langevin. A sample of La1.855Sr0.145CuO4 crystal, consisting of two crystals (total mass 3.5 g, = 36 K), was cut from a single rod grown via the traveling-solvent floating-zone method and co-aligned within less than . The sample with lattice parameters Å, Å in the high-temperature tetragonal unit cell notation, was measured in a standard cryostat. Magnetic excitation spectra were measured with a fixed sample orientation (axis perpendicular to the horizontal scattering plane) for 12 or 24 hours at different temperatures. Two different setups were employed, with incident wavelengths and 5.0 Å (see Table 1). The shorter wavelength was used to confirm the size of the superconducting spin gap (4 meV)KofuPRL2009; ChangPRL2009 at the base temperature of = 2 K (see Supplementary Material (SM) Note 1). The = 5.0 Å setup was used to collect spectra at = 40, 30, 25 and 20 K with a resolution of = 80 eV.
| Setup | (Å) | (eV) | interval (meV) |
| I | 3.3 | 384 | and |
| II | 5.0 | 80 | and |
III Results
The neutron scattering rate is proportional to the four-dimensional dynamic structure factor , which is directly connected to the dynamic magnetic susceptibility: with a proportionality factor equal to the Bose occupation factor = 1/[exp()-1] for neutron energy-gain processes () and to for energy-loss processes (). Here, represents the three dimensional momentum transfer, the energy transfer and the Boltzmann constant. As the electronic band structure and spin excitations of La2-xSrxCuO4 are quasi two-dimensional HorioPRL2018; RoemerPRB2015; Lake2005, momentum transfers (,,) were projected to (,,0)ChristensenPRL2004; VignolleNATPHYS2007; MonneyPRB2016. Figure 2a-d illustrates the resulting scattered neutron intensity of La1.855Sr0.145CuO4 in the -plane, integrated over an energy range of = [-2.15,-0.15] meV for = 40, 30, 25 and 20 K. A step size of 0.0075 r.l.u. was chosen for this plot. The -map at = 40 K reveals an enhanced neutron intensity at wavevectors =(,1/2,0) and =(,,0) with . These excitations, known to arise from the magnetic interactions in La2-xSrxCuO4 Dean2013, appear as straight columns in the measured energy range because the spin-wave bandwidth amounts to300 meV. Their reciprocal lattice positions coincide with the positions where magnetic Bragg peaks appear at slightly lower Sr doping or under magnetic field ChangPRB2008; ChangPRL2009. Thus, we interpret the excitations as signatures of the nearby ordered state.
When the temperature is reduced below the superconducting transition temperature ( = 36 K) the magnetic excitation spectrum is gradually modified (see Fig. 2). Detailed insight into the temperature dependence of the low-energy fluctuations is gained in Fig. 2e-h, where the energy dependence of the magnetic fluctuations is shown along (,0.5,0). Here, we used an integration width along (0, 0.5 0.025, 0) r.l.u, a step sizes along and 60 eV along the energy axis. The shaded area extends over one and two elastic FWHMs, colored in dark and light gray, respectively. The data at = 40 K shows two excitation branches at =(,1/2,0) with energy dependent intensity modulation peaking below an energy transfer of 1 meV. Upon cooling below the magnetic intensity is gradually suppressed, but our results suggest that magnetic fluctuations persist at low energy transfers and remain visible down to 20 K (c.f. Fig. 2g).
An analysis of this behavior was carried out through -cuts at constant energy transfers, using an optimized integration width (0.025 r.l.u.) and step size to obtain the best signal-to-background ratio (see SM Note 2). Data at various energy transfers were integrated over 0.1 meV intervals. Cuts along (,1/2,0) and (1/2,,0) for different and temperatures were fitted with two Gaussians on a constant background. Since no significant variances between cuts along and were observed (see SM Note 3), we combined and scans to produce Fig. 3. We furthermore used detailed balance, = exp() ), to combine energy-loss and energy-gain spectra (see SM Note 4). Figure 3 displays the dynamic structure factor for different temperatures across the excitation branches at various excitation energies. We note that the -cuts have been offset for clarity. At = 40 K, two Gaussian peaks at =(,1/2,0) are observed over the entire probed energy range. The magnetic fluctuations are significantly weakened below = 36 K but remain visible up to 2 meV at = 30 K. Deeper inside the superconducting phase at = 25 K, the two excitation branches are observed below 1 meV. These two excitations are largely suppressed above 0.35 meV at = 20 K.
This non-uniform energy dependence of the magnetic fluctuations is shown in Fig. 4 for the different temperatures. Here, the data at =(,1/2,0) and =(,,0) were combined. At each excitation energy and temperature, the -cuts were fitted to Gaussian lineshapes, and the average amplitude at the peak center is represented as one point in Fig. 4 (see SM Note 5 for details). The subpanels compare pair-wise for = 40, 30, 25 and 20 K. They confirm the behavior observed in Fig. 3 that low-energy magnetic fluctuations persist to low temperature and that the overall magnetic spectral weight is suppressed (see also SM Note 6). As guides to the eyes the higher temperature of each subpanel was fitted to an effective spectral lineshape consisting of four Gaussians on a constant background (see solid lines in Fig. 4), where the two Gaussians on the neutron energy-loss side were constrained to the ones on the energy-gain side using detailed balance. The dashed lines in the panels simulate the expected variation of the solid line in a scenario when the temperature dependence is attributed to the Bose occupation factor only. This unambiguously demonstrates that the observed suppression of the low-energy magnetic fluctuations at = 30, 25 and 20 K does not arise alone from thermal depopulation, but is connected to the another effect impacting the dynamic susceptibility.
IV Discussion
The opening of a superconducting spin gap provides a natural explanation for the suppression of low-energy fluctuations ChangPRL2009; KofuPRL2009; LakeNAT2002. However, by analogy to a gap in the electronic spectrum one would expect that, as the sample is cooled below the spin gap should initially be detected at the lowest energy transfers, and that higher-energy excitations should only be wiped out at lower temperatures. Instead, we observe that the magnetic fluctuations are suppressed equally over an energy window of 3 meV (see SM Note 6 for details), and that low energy fluctuations below 1 meV persist over an appreciable temperature window down to roughly 20 K. Thus, our results indicate the presence of enhanced low-energy spectral weight in La1.855Sr0.145CuO4, as indicated by the Gaussian fits in Fig. 4.
In a study KofuPRL2009 of the low-energy fluctuations in La2-xSrxCuO4 over the composition-range a two-component spectral response bearing similarity to our observations was observed at base temperature in samples displaying magnetic order. Here, it was hypothezised that stripe-ordered domains with associated low-energy glassy fluctuations coexist with superconducting regions, displaying a spin gap of order KofuPRL2009. Following this interpretation, the low-energy components of the temperature-dependent spectrum observed in La1.855Sr0.145CuO4 correspond to critical fluctuations of the incipient stripe-ordered phase. Because La1.855Sr0.145CuO4 resides on the disordered side of the stripe quantum critical point these slow fluctuations eventually succumb to the opening of the spin gap, while they survive in the magnetically ordered samples studied in Ref. KofuPRL2009.
The interpretation of the persistent low-energy fluctuations as the harbingers of a proximate ordered state, at first sight appears at odds with the prior observation by some of us that the application of a magnetic field perpendicular to the CuO2 planes results in a gradual decrease of the spin gap, and that magnetic order appears when this gap closes ChangPRL2009. We remark, however, that vortices may not only suppress superconductivity and with it the spin gap, but could also facilitate a slowing down of the low-energy critical fluctations. This could have been missed in Ref. ChangPRL2009 because the energy resolution was significantly worse than what we obtained in the present experiment. Thus, further high-resolution experiments on other La2-xSrxCuO4 compositions will be fruitful to check this hypothesis, and to clarify the nature of the low-energy fluctuations below 1 meV.
In a more general context we notice that cases with two coexistent static order parameters displaying competing tendencies are usually referred to as phase competition. Theoretically such systems are often described by a single complex-valued order parameter BrussPRB1999; Zhang1997, where the competing interaction provides an adiabatic continuity between the two phases (see Fig. 1a). Our study suggests that the phase competition can be extended beyond the static co-existence phase. When only one phase has condensed into the ground state (in this case superconductivity), a dynamic phase competition indicates a state where the fluctuations of the other order parameter shows competing tendencies. La1.855Sr0.145CuO4 features at least three non-thermal routes through extrinsic tuning to static AFM order KhaykovichPRB2005; ChangPRB2008; ChangPRL2009; LakeNAT2002. Thus, the AFM paramagnetic fluctuations are about to condense at K, just above the superconducting transition. As the superconducting gap opens, particle-hole excitations are suppressed, leading to a competition between the condensation of unconventional Cooper-pairs and AFM fluctuations. In La1.855Sr0.145CuO4 superconductivity dominates this competition, but extrinsic tuning shows an adiabatic connection to static magnetic order where the magnetic modes soften into the ground state. Notably, the low energy scale of the observed enhanced spectral weight below 1 meV may explain why only small perturbations (Sr concentration 0.135 and 7 T) induces magnetic long-range order.
The concept of a dynamic competition between superconductivity and magnetism is likely existing also in other materials. In the heavy-fermion materials such as CeCu2Si2 for instance, magnetism and superconductivity possess very similar onset temperatures StockertNATPHYS2011. The spin gap (0.3 meV) is, however, an order of magnitude smaller than in La1.855Sr0.145CuO4 StockertNATPHYS2011, which in turn suggests a dynamic competition on a much smaller energy scale. For many pnictides, AFM order sets in above superconductivity, leading to a static competition between magnetism and superconductivity NandiPRL2010. However, close to the magnetic quantum critical point, dynamic competition should arise. For instance, in Ba0.6K0.4Fe2As2, the spin resonance is sufficiently high in excitation energy enabling future investigations of dynamic competition between magnetism and superconductivity with neutron spectroscopy ChristiansonNAT2008.
V Outlook
Finally, it is interesting to consider dynamic competitions between quantum phases beyond the interplay between magnetism and superconductivity. The field of quantum materials features a plethora of systems with competing phases, but in many cases a dynamic competition is avoided. In some cases, associated fluctuations manifest at different wavevectors. This is the case in CeCu5.8Ag0.2, for instance Poudel2019. In this material spin-density wave-like fluctuations condensate into the ground state establishing magnetic order at higher Ag concentration, while the competing fluctuations remain gapped. The superconducting ground state is special in this regard, because the Cooper pair condensation leads to a gap over an appreciable range in reciprocal space which easily impacts other fluctuations. Thus, dynamic competition may be expected also in the interplay between superconductivity and charge order, and materials featuring spin- and charge-density fluctuations. Future experiments could be beneficial to better understand the concept of dynamic competition between different quantum phases, and compare it to cases - such as frustrated magnetic materials - where a a dynamic cooperation among low-energy fluctuations is established Balents2010; Broholm2020.
Acknowlegdements: The time-of-flight neutron experiments were performed at the spectrometer IN5 Ollivier2011 of the Institut Laue-Langevin, France. The experimental data used in this work can be found using the link https://doi.org/10.5281/zenodo.10992720. We acknowledge beam line support and technical advices from H. Mutka. This work was supported by the Swiss National Science Fundation (through NCCR, MaNEP, and grant Nr PZ00P2-142434) and the Ministry of Education and Science of Japan.
Supplementary Material for: Dynamic Competition between Cooper-Pair and Spin-Density-Wave Condensation
B. Decrausaz1,2, M. Pikulski3, O. Ivashko1, N.B. Christensen4, J. Choi5, L. Udby6,
Ch. Niedermayer2, K. Lefmann6, H.M. Rønnow7, J. Mesot2,3, J. Ollivier8, T. Kurosawa9,10,
N. Momono9,10, M. Oda9, J. Chang1111Email: johan.chang@physik.uzh.ch, D.G. Mazzone2222Email: daniel.mazzone@psi.ch
1Physik-Institut, Universität Zürich, Winterthurerstrasse 190, CH-8057 Zürich, Switzerland
2PSI Center for Neutron and Muon Sciences, 5232 Villigen PSI, Switzerland
3Laboratory for Solid State Physics, ETH Zürich, 8093 Zürich, Switzerland
4Department of Physics, Technical University of Denmark, DK-2800 Kongens Lyngby, Denmark
5Diamond Light Source, Harwell Campus, Didcot OX11 0DE, United Kingdom
6Nano-Science Center, Niels Bohr Institute, University of Copenhagen, DK-2100 Copenhagen, Denmark
7Laboratory for Quantum Magnetism, EPFL, CH-1015 Lausanne, Switzerland
8Institut Laue-Langevin, BP 156, F-38042 Grenoble, France
9Department of Physics, Hokkaido University, Sapporo 060-0810, Japan
10Department of Applied Sciences, Muroran Institute of Technology, Muroran 050-8585, Japan
(Dated: September 25, 2025)
Supplementary Note 1. Confirmation of the Superconducting Spin Gap
A setup with an incoming neutron wavelength of Å was used to confirm the size of the superconducting spin gap. The resulting intensity is depicted in Fig. 1 as function of (, 0.5, 0) in r.l.u and excitation energy for various temperatures Fig. 2 presents one-dimensional cuts through the data along = (, 0.5, 0) for different energies. Our observations are consistent with 4 meV found previouslyKofuPRL2009; ChangPRL2009. In Fig. 1a and Fig. 2a, we show the spin excitations in the normal state, at = 40 K . Incommensurate excitations are observed down to the lowest measurable excitation energies. At base temperature = 2 K the intensity was to low to extract the gap size because of the Bose factor reduction (see Fig. 2e). However, at = 10 K the spectral weight up to 4 meV is gapped out. The enhanced intensity near 2 meV in Fig. 1 is attributed to parasitic scattering from the IN5 sample environment.
Supplementary Note 2. Optimized Integration Width in
The four dimensional data obtained from IN5 was reduced to one dimension, integrating over the other three dimensions. Along the data was integrated over the full range covered by the data set as the spin excitations in LSCO are assumed to be two dimensional tranquadaNeutronscatteringStudyStripephase1996; lakeThreedimensionalityFieldinducedMagnetism2005. To assess the optimal integration width along and we first investigated various fixed energy integration widths at different temperatures (see Table 2). They were chosen to assess whether a single integration width can be used over the entire energy range. We note that the 20 K data were neglected during this procedure due to the weak signal.
-cuts with varying integration widths between 0.01 and 0.1 (r.l.u.) were generated around the four spin-density wave wavevectors, which we labeled peak 1 to 4 for =(,1/2,0) and =(,,0), respectively (as examples of -cuts see Fig. 3a and Fig. 3b). They were fitted with two Gaussian lineshapes on a constant background. We calculated the ratio between the amplitude and background (A/B) for each peak and plotted them against the integration width (see Fig. 4, 5, 6). The optimal integration width in was chosen to the value where the A/B ratio no longer improves. We also fitted one of the resulting peak pairs of the -cuts and -cuts to a Gaussian on a sloping background. These fits were overlayed with the results of the other two peaks (see Fig. 4 for instance). From these results we concluded that identical integration widths for a particular -direction can be used. The overall results are summarized in Table 3, showing that an integration width of 0.025 (r.l.u.) is ideal for both and cuts at all temperatures.
| Temperature | 40 K | 30 K | 25 K |
| energy integration width in (meV) | [-1.9, -0.9] | [0.5, 1.5] | [-0.75, -0.15] |
| Temperatures | cut width of peak 1 and 2 | cut width of peak 3 and 4 |
| 40 K | 0.033 0.007 (r.l.u.) | 0.026 0.026 (r.l.u.) |
| 30 K | 0.013 0.036 (r.l.u.) | 0.023 0.023 (r.l.u.) |
| 25 K | 0.05 0.005 (r.l.u.) | 0.023 0.015 (r.l.u.) |
Supplementary Note 3. Combination of - and -cuts
At 40 K, we compared the center and the full-width at half maximum (FWHM) of all , -cuts – see supplementary Fig. 7 and Fig. 8. From this, we conclude that the peaks are statistically indistinguishable. On this basis, we have folded the and scans into single branches. This operation improves counting statistics while retaining the peak position and peak width. In our analysis, we have applied this operation to all temperatures.
Supplementary Note 4. Detailed Balance
An assessment of the -cuts at negative and positive energy transfers revealed that the data can be further reduced using detailed balance: = with = exp(-/). Figure 9 shows various -cuts at different energy transfers and temperatures. For the cuts where a corresponding energy on the energy loss side was measured both cuts are overlayed, the energy gain spectrum and the corrected energy loss cuts. We note that we needed to further correct for a constant background, stemming form background contributions attributed to the sample environment. Notably, detailed balance corrects for the temperature of the sample but not the one of the cryostat usually remaining around 200 K.
Supplementary Note 5. Amplitude vs. Energy Plots
The spectral weight of the spin excitations was investigated at the fitted -positions of the combined - and (see Supplementary Note 3) Figure 10 shows the spectral weight at = 40, 30, 25 and 20 K as a function of energy using an energy step size of 0.1 meV, energy integration width of 0.1 meV in the range of = -3.15 to = -0.15 meV and = 0.15 to = 1.65 meV (omitting the elastic line). was integrated over = [0.4875, 0.5125] rlu. The individual points of these figures result from a two step fitting process. Firstly, individual -cuts were fitted with two Gaussians. The weighted average of the resulting FWHM and center of the 40 K data were then used as fixed parameters to refine the amplitudes, background and slope of the two Gaussians for the data at all temperatures. Figure 10 shows the obtained amplitudes as function of energy at different temperatures. Since the spectral weight of both reduced peaks match, we concluded that the spectral weight of all four excitation branches can be folded onto a single branch.
Supplementary Note 6. Temperature dependent suppression of low-energy spectral weight
In this section we add supplementary material for the temperature dependent suppression of the low-energy spectral weight. Figure 11a shows a temperature difference plot of the neutron intensity along the (, 0.5, 0) and energy-transfer axes. The plot results from the data shown in Fig. 2 of the main manuscript, for which the Bose-corrected = 40, 30 and 25 K datasets were subtracted with the Bose-corrected = 20 K dataset. In panel (b) we report on the relative suppression of the spectral weight at (,) as a function of energy transfer. Here the energy dependence at = 30, 25 and 20 K shown in Fig. 4 of the main manuscript has been normalized by the = 40 K dataset. Within the employed counting statistics we find an equivalent suppression of the low-energy spectral weight across all energy transfers. Notably, the mean spectral weight is suppressed by 60, 80 and 90 % when the = 35, 25 and 20 K data are compared to the = 40 K data. In panel (c) we show the temperature dependence of (,) at specific energy transfers, corroborating the finding of panel (b). We note that the chosen match the ones of the Q-cuts in Fig. 3 of the main manuscript.